Dr. Bryan Carpenter

Department of Electronics and Computer Science

August 1, 1995

Computing Services University of Southampton

Dear Computing Services,

I am enclosing a request for use of the UNIX workstations in building 44 for a tutorial which we are holding on October 10th. This is part of a two day workshop on High Performance Fortran.

For this tutorial we will need to install our High Performance Fortran translation system on the local servers. As well as the translator itself, this requires the following software to be installed

• A C++ compiler.

• A Fortran 90 compiler.

• An MPI (Message Passing Interface) installation.

The favoured combination is NAG f90, GNU C++ and Argonne MPICH.

David Adams tells me that NAG Fortran 90 can be made visible on these machines if we obtain permission from NAG, which we expect to be able to do. MPICH is public domain, and can easily be installed (but I understand from Ian Hardy that there are some restrictions on remote shell execution on those machines, which may cause problems).

The exact amount of disk space needed will depend on which of these systems are already available on the workstations.

Thank you.